# Trackable Agent-based Evolution Models at Wafer Scale

### v2

@@ Introduction
Take, for instance, the Avida platform, which instantiates populations of self-replicating computer programs for evolution experiments. When running on a single CPU, this system can support about 20,000 generations per day, or about two hundred million individual replication cycles daily (ofria2009artificial). By way of comparison, E. coli populations within individual flasks of the Lenski Long-Term Evolution Experiment undergo only six doublings per day, meaning their generations take orders of magnitude longer than Avidians (good2017dynamics). (In continuous culture, though, the rate can be c. 72 generations per day.) Indeed, such capability for fast generational turnover has been a key motivation for using artificial life systems to study evolution. However, the effective population size of flasks in the Long-Term Evolution Experiment is orders of magnitude larger than Avida’s: 30 million vs. 10,000. Consequently, these systems actually exhibit a similar number of replication events per day. This pattern of dramatically faster generation times than those observed in nature and dramatically smaller populations largely generalizes across artificial life systems. Of course, any such comparisons should also note profound discrepancies between the genetic, phenotypic, and environmental richness of biological organisms and alife models.
Crucially, however, the scale of population size can greatly impact subjects of artificial life research, like transitions in individuality, ecological dynamics, and rare evolutionary innovations (taylor2016open; dolson2021digital; taylor2019evolutionary). Cross-scale dynamics are also crucial to many key real-world challenges. For example, in evolutionary epidemiology, interactions between within-host infection dynamics and population-level epidemiological patterns determine the evolutionary trajectory of the population (schreiber2021cross). However, because capabilities of current silicon-based processors are not expected to improve markedly in the foreseeable future (sutter2005free), scale-up necessary to progress on these key frontiers will demand many-processor computation. Application of parallel and distributed computation, however, imposes compromises to the convenience, flexibility, observability, interpretability, total reliability, and perfect replicability enjoyed under classical centralized, serial models of computation. Encouragingly, these challenges are already implicit to much of biology; the productivity of research involving natural organisms evidences that they are surmountable and even hints at strategies that can be used to solve them. Here, we explore alignment of digital evolution to HPC accelerator hardware at the extreme cutting edge of massively distributed computation, and use techniques inspired by those applied to natural organisms to mitigate limitations of distributed computation with respect to tracking phylogenies.
@@
Achieving highly scalable artificial life and digital evolution systems involves two distinct engineering considerations. First, as with any high-performance scientific computing, careful design is required to appropriately divvy computation and communication workloads across available hardware. Second, given the exceptionally discretionary nature of artificial life modeling, we can intentionally tailor simulation semantics to suit underlying hardware capabilities. Ackley’s ongoing work with the T2 Tile Project and ULAM exemplifies a strong synthesis of this engineering duality (ackley2016ulam). At the level of simulation semantics, Ackley formulates update procedures in terms of local interactions between discrete, spatially situated particles. This design provides for efficient one-to-one mapping between simulation space and hardware components, minimizing requirements for intra-hardware connectivity and preventing global impacts from on-the-fly augmentations or reductions of available hardware. The ULAM framework then ties into implementation-level infrastructure necessary to accomplish performant, best-effort lock/release of spatial event windows spanning bordering hardware unites (ackley2013movable). Ackley’s work is distinguished, in fact, in dipping to a yet-lower level of abstraction and tackling design of bespoke, modular distributed processing hardware (ackley2011homeostatic; ackley2023robust; livingcomputationSFBSanta).
@@
A major upshot of the deep learning race is the emergence of spectacularly capable next-generation compute accelerators (zhang2016cambricon; emani2021accelerating; jia2019dissecting; medina2020habana). Although tailored expressly to deep learning workloads, these hardware platforms represent an exceptional opportunity to leapfrog progress on grand challenges in artificial life. The emerging class of fabric-based accelerators, led by the 850,000 core Cerebras CS-2 Wafer Scale Engine (WSE) (lauterbach2021path; lie2022cerebras), holds particular promise as a vehicle for artificial life models. This architecture interfaces multitudinous processor elements (PEs) in a physical lattice, with PEs executing independently with private on-chip memory and interacting locally through a network-like interface.
Our work here explores how such hardware might be recruited for large-scale digital evolution, demonstrating a genetic algorithm implementation tailored to the dataflow-oriented computing model of the CS-2 platform. Indeed, rapid advances in the capability of accelerator devices, driven in particular by market demand for deep learning operations, are anticipated to play a key role in future advances in agent-based model capabilities (perumalla2022computer). The very recent CS-3 chip announcement, which boosts clustering support to potentially thousands of constituent accelerators (moore2024cerebras).
@@
Unfortunately, scale complicates simulation observability. It is not uncommon for the volume and velocity of data streams from contemporary simulation to outstrip hardware bandwidth and storage capacity (osti_1770192). Intentional engineering effort will be required to ensure large-scale simulation retains utility in pursuing rigorous hypothesis-driven objectives.
Here, we confront just a single aspect of simulation observability within distributed evolutionary simulation: phylogenetic history (i.e., evolutionary ancestry relationships). Phylogenetic history plays a critical role in many evolution studies, across study domains and in vivo and in silico model systems alike (faithConservationEvaluationPhylogenetic1992; STAMATAKIS2005phylogenetics; frenchHostPhylogenyShapes2023; kim2006discovery; lewinsohnStatedependentEvolutionaryModels2023a; lenski2003evolutionary). Phylogenetic analysis can trace the history of notable evolutionary events (e.g., extinctions, evolutionary innovations), but also characterize more general questions about the underlying mode and tempo of evolution (moreno2023toward; hernandez2022can; shahbandegan2022untangling; lewinsohnStatedependentEvolutionaryModels2023a). Particularly notable, recent work has used comparison of observed phylogenies against those produced under differing simulation conditions to test hypotheses describing underlying dynamics within real-life evo-epidemiological systems (giardina2017inference; voznica2022deep). Additionally, in silico, phylogenetic information can even serve as a mechanism to guide evolution in application-oriented domains (lalejini2024phylogeny; lalejini2024runtime; murphy2008simple; burke2003increased).
@@
In order to analyze the phylogenetic history of an evolving system, we must first collect it. In existing artificial life work, a typical approach uses centralized tracking to maintain an exact, complete record comprising all parent-offspring relationships that have existed over the course of a simulation (ray1992evolution; bohm2017mabe; de2012deap; garwood2019revosim; godin2019apoget; dolson2024phylotrackpy). Often, records of extinct lineages are pruned to prevent memory bloat (moreno2024analysis). Although the direct-tracking approach is well suited to serial simulation or centralized controller-worker schemes, extreme sensitivity to missing data and runtime communication overheads impede scaling to highly distributed systems — particularly those with lean memory capacity like the Cerebras WSE (moreno2024analysis). To overcome this limitation, we have developed reconstruction-based approaches to phylogenetic tracking in silico (moreno2022hereditary). These approaches require no centralized data collection during simulation runtime; instead, they use post hoc comparisons among end-state agent genomes to deduce approximate phylogenetic history — akin to how DNA-based analyses tell how lineages of contemporary organisms relate in natural history. Figure 1 summarizes this reconstruction-based strategy.
Although analogous work with natural biosequences is notoriously challenging and data-intensive (neyman1971molecular; lemmon2013high), recently-developed hereditary stratigraphy annotation architecture is explicitly designed for fast, accurate, and data-lean reconstruction. Designed for attachment to underlying replicators as a neutral annotation (akin to noncoding DNA), it is a general-purpose technique potentially applicable across study domains (liben2008tracing; cohen1987computer; friggeri2014rumor). In a stroke of convergent thinking, Ackley’s recent work reports use of “bar code” annotations on his self-replicators to achieve a measure of coarse-grained lineage tracing (ackley2023robust).
@@
In this paper, we report new software and algorithms that harness the Cerebras Wafer Scale Engine to enable radically scaled-up agent-based evolution while retaining key aspects of observability necessary to support hypothesis-driven computational experiments. Implementation comprises two primary aspects:
@@ Methods
This section details the algorithm and implementation design used in this work. After briefly recounting the core mechanics of hereditary stratigraphy methods for phylogenetic tracking, we motivate and describe new lightweight “surface” data structures used to annotate agent genomes in this work. Then, we describe the asynchronous compute-communicate implementation strategy behind the testbed WSE-oriented island-based genetic algorithm used for validation and benchmarking experiments.
@@
In biology, mutational drift encodes ancestry information in DNA genomes. Our proposed methods are analogous, with ancestry information captured within genomes themselves rather than external tracking. Phylogenetic history can then be estimated after the fact, as Figure 1 depicts. This strategy reduces runtime communication and is resilient to germane modes of data loss (e.g., dropped messages, hardware crash-out).
@@
Small differentia size intensifies need for a lean data structure to back differentia record management. Shrinking differentia to a single bit would be absurd if each is accompanied by a 32-bit generational timestamp. To safely drop timestamps, though, we need means to recalculate them on demand. As such, all described algorithms include means to deduce timestamps of retained differentia solely from their storage index and the count of the record’s elapsed generations.
Lastly, design of hstrat annotations must also consider how available storage space should be allocated across the span of history. In one possible strategy, retained time points would be distributed evenly across history. In another, more recent time points would be preferred. We term these as “steady” and “tilted” policies, respectively (han2005stream; zhao2005generalized). Note that prior hereditary stratigraphy work refers to them instead as “depth-proportional” and “recency-proportional resolution” (moreno2022hereditary). Comparisons of reconstruction quality have shown that tilted policy gives higher quality reconstructions from the same amount of reconstruction space in most — but not all — circumstances (moreno2024guide). This pattern follows an intuition that high absolute precision is more useful to resolving recent events than more ancient ones. In practice, it may be desirable to use a hybrid approach that allocates half of available annotation space to each strategy (moreno2024guide). The bottom panels of Figure 3 contrast steady versus tilted behaviors.
@@
At the outset of this project, several problematic aspects of porting existing hereditary stratigraphy algorithms to the WSE became apparent. Issues stemmed, in part, from a fundamental feature of these algorithms: organization of retained strata in contiguous, sorted order. Forthwith, we refer to algorithms following this convention as “column”-based. The column design, and attendant implementation details, impose various drawbacks:
@@
Addressing these concerns required fundamental reformulation of hereditary stratigraphy conventions. To this end, we introduce a constant-time indexing scheme that maps each lineage’s fingerprint stream directly onto a fixed-width memory buffer. Differentia pruning occurs implicitly, as a result of resident differentia being overwritten by new placements. As before, care is taken to guarantee temporally representative collections of resident differentiae within the memory buffer over elapsed time. We refer to this new approach as “surface”-based, as opposed to “column”-based. Figure 2 contrasts the two.
steady retention policy
tilted retention policy
@@
The Cerebras Wafer-Scale Engine comprises a networked grid of independently executing compute cores (Processing Elements or PEs). Each PE contains special message-handling infrastructure, which routes tagged 32-bit packets (“wavelets”) to neighboring PEs and/or to be processed locally, according to a programmed rule set. Each Processor Element is equipped with 48kb of private on-chip memory, which can be accessed within a single clock cycle; communication to neighboring PEs, too, incurs low latency (buitrago2021neocortex). Processor Elements provide faculties for standard arithmetic and flow-control operations, as well as vectorized 32- and 16-bit integer and floating-point operations.
The WSE device is programmed by providing “kernel” code, written in the Cerebras Software Language (CSL), to be executed on each PE. This language’s programming model purposefully reflects underlying capabilities and particularities of the WSE architecture. CSL organizes code within an event-driven framework, with the programmer defining tasks to be triggered in response to wavelet activation signals exchanged between — and within — PEs. Scheduling of active tasks occurs via hardware-level microthreading, which allows for some level of concurrency. Special faculties are provided for asynchronous send-receive operations that exchange contiguous, or strided, buffer data between PEs.
@@
Our implementation unfolds according to a generational update cycle, schematized in Figure 4. Migration, depicted as blue-red arrows, is handled first. Each PE maintains independent immigration buffers and emigration buffers dedicated to each cardinal neighbor, depicted by Figure 4 in solid blue and red, respectively. On simulation startup, asynchronous receive operations are opened to accept genomes from each neighboring PE into its corresponding immigration buffer. At startup, additionally, each emigration buffer is populated with genomes copied from the population and an asynchronous send request is opened for each. Asynchronous operations are registered to on-completion callbacks that set a per-buffer “send-” or “receive-complete” flag variable. In this work, we size send buffers to hold one genome and receive buffers to hold four genomes. The main population buffer held 32 genomes.
Subsequently, the main update loop tests all completion flags. For each immigration flag that is set, buffered genomes are copied into the main population buffer, replacing randomly chosen population members. Then, the flag is reset and a new receive request is initiated. Likewise, for each emigration flag set, corresponding send buffers are re-populated with randomly sampled genomes from the main population buffer. Corresponding flags are then reset and new send requests are initiated. The bottom right corner of Figure 4 summarizes the interplay between send/receive requests, callbacks, flags and buffers.
@@
Genomes were handled as fixed-length segments of raw memory. We used a 96-bit genome for clade reconstruction trials, shown in Supplementary Figure 8 (moreno2024supplement). At the outset of simulation, the first 16 bits of founding genomes were randomized and, subsequently, were inherited without mutation, thus identifying descendants of the same founding ancestor. The next 80 bits were used for hereditary stratigraph annotation, 16 bits for a generation counter and the remaining 64 as a field for single-bit differentia managed according to a tilted surface policy. For other experiments, a 128-bit genome layout was used instead, with the first 32 bits used for a floating point “fitness” value and the generation counter upgraded from 16 to 32 bits.
@@
Python implementations of the surface tilted and steady algorithms both took around 4.2 microseconds per operation (SEM approx. 50; ). For context, this speed was about the measured time for a surface placement using a trivial calculation (SEM approx. 0.05; ). As expected, column implementations of steady and tilted fared much worse, taking about and the execution time per operation compared to the surface operations. In both cases, surface implementations significantly outperformed their column counterparts (Mann-Whitney U test, ). It should be noted that these results reflect an intrinsic performance penalty due to interpreter overhead of Python evaluation. However, comparisons among these Python implementations are nonetheless informative because all are on equal footing in this regard.
Zig microbenchmarks clock tilted surface annotation updates at 233 ns per operation (SEM 0.9; ). For context, this time is a little more than twice that required for a main memory access in contemporary computing hardware (markus2022memory). Our results measure the operation at the measured time for a trivial placement calculation (SEM 0.2; ).
@@
Multiplied out to a full day, 17 generations per second turnover would elapse around 1.5 billion generations. With 32 individuals hosted per each of 850,000 PEs, the net population size would sit around 27 million at full CS-2 wafer scale. (Note, though, that available on-chip memory could support thousands of our very simple agents per PE, raising the potential for a net population size on the order of a billion agents.) A naive extrapolation estimates on the order of a quadrillion agent replications could be achieved hourly at full wafer scale for such a very simple model. We look forward to conducting more thorough benchmarking and scaling experiments in future work.
@@
These timings indicate cost of phylogeny tracking is approximately equivalent to that of the other aspects of simulation, combined. Given the highly minimalistic nature of the agent model and selection process, this result is highly promising. In actual use, most experiments will likely involve a much more sophisticated agent model; thus, the relative overhead of tracking will diminish significantly. It is additionally worthwhile to note that caching and coarsening strategies discussed above have not yet been applied, which may reduce the performance penalty of tracking considerably.
@@
Here, we demonstrate phylogenetic reconstructions from emulated WSE hardware and inspect clade structure to qualitatively assess their credibility. For these experiments, we tagged genomes with a 16-bit randomly generated identifier at simulation startup, as shown in Supplementary Figure 8 (moreno2024supplement). Using the hardware emulator, we instantiated populations over a PE collective for durations of 25, 50, 100, and 250 generation cycles under drift conditions. At the end of simulation, phylogenetic reconstruction was conducted using one genome sampled per PE.
Figure 6 shows reconstructed phylogenies from each duration. As expected, taxa belonging to the same founding clade (shown by color) are reconstructed as more closely related to each other than to other taxa. Additionally, and also as expected, the number of distinct remaining founding clades diminishes monotonically with increasing simulation duration. These consistencies corroborate the general integrity of implemented annotation and reconstruction processes. Note, however, the incidence of moderate overestimations for relatedness between independently-tagged clades throughout. As mentioned earlier, and discussed in greater depth elsewhere (moreno2024guide), this is an expected artifact of hereditary stratigraphy with single-bit differentia. Applications requiring greater reconstruction precision can opt for larger differentia sizes and/or higher differentia counts.
@@
Finally, we set out to assess the performance of our pipeline at full wafer scale. For these experiments, we used the four-word, tracking-enabled genome layout, with the full first 32 bits containing a floating point fitness value. We defined two treatments: purifying-only, where 33% of agent replication events decreased fitness by a normally-distributed amount, and adaption-enabled, which added beneficial mutations that increased fitness by a normally-distributed amount, occurring with 0.3% probability. These beneficial mutations introduced the possibility for selective sweeps. As before, we used tournament size 5 for both treatments. We performed four on-hardware replicates of each treatment instantiated on 10k (), 250k () and 562.5k () PE arrays. We halted each PE after it elapsed 1 million generations.

### v3

@@ Introduction
Take, for instance, the Avida platform, which instantiates populations of self-replicating computer programs for evolution experiments. When running on a single CPU, this system can support about 20,000 generations per day, or about two hundred million individual replication cycles daily (ofria2009artificial). By way of comparison, E. coli populations within individual flasks of the Lenski Long-Term Evolution Experiment undergo only six doublings per day, meaning their generations take orders of magnitude longer than Avidians (good2017dynamics). (In continuous culture, though, the rate can be c. 72 generations per day.) Indeed, such capability for fast generational turnover has been a key motivation for using artificial life systems to study evolution. However, the effective population size of flasks in the Long-Term Evolution Experiment is orders of magnitude larger than Avida’s population size: 30 million vs. 10,000. Consequently, these systems actually exhibit a similar number of replication events per day. This pattern of dramatically faster generation times than those observed in nature and dramatically smaller populations largely generalizes across artificial life systems. Of course, any such comparisons should also note profound discrepancies between the genetic, phenotypic, and environmental richness of biological organisms and ALife models.
Crucially, however, the scale of population size can greatly impact subjects of artificial life research, like transitions in individuality, ecological dynamics, and rare evolutionary innovations (taylor2016open; dolson2021digital; taylor2019evolutionary). Cross-scale dynamics are also crucial to many key real-world challenges. For example, in evolutionary epidemiology, interactions between within-host infection dynamics and population-level epidemiological patterns determine the evolutionary trajectory of the population (schreiber2021cross). However, because capabilities of current silicon-based processors are not expected to improve markedly in the foreseeable future (sutter2005free), the scale-up necessary to progress on these key frontiers will demand many-processor computation. Application of parallel and distributed computation, however, imposes compromises to the convenience, flexibility, observability, interpretability, total reliability, and perfect replicability enjoyed under classical centralized, serial models of computation. Encouragingly, these challenges are already implicit to much of biology; the productivity of research involving natural organisms evidences that they are surmountable and even hints at strategies that can be used to solve them. Here, we explore alignment of digital evolution to HPC accelerator hardware at the extreme cutting edge of massively distributed computation, and use techniques inspired by those applied to natural organisms to mitigate limitations of distributed computation with respect to tracking phylogenies.
@@
Achieving highly scalable artificial life and digital evolution systems involves two distinct engineering considerations. First, as with any high-performance scientific computing, careful design is required to appropriately divvy computation and communication workloads across available hardware. Second, given the exceptionally discretionary nature of artificial life modeling, we can intentionally tailor simulation semantics to suit underlying hardware capabilities. Ackley’s ongoing work with the T2 Tile Project and ULAM exemplifies a strong synthesis of this engineering duality (ackley2016ulam). At the level of simulation semantics, Ackley formulates update procedures in terms of local interactions between discrete, spatially situated particles. This design provides for efficient one-to-one mapping between simulation space and hardware components, minimizing requirements for intra-hardware connectivity and preventing global impacts from on-the-fly augmentations or reductions of available hardware. The ULAM framework then ties into implementation-level infrastructure necessary to accomplish performant, best-effort lock/release of spatial event windows spanning bordering hardware units (ackley2013movable). Ackley’s work is distinguished, in fact, in dipping to a yet-lower level of abstraction and tackling design of bespoke, modular distributed processing hardware (ackley2011homeostatic; ackley2023robust; livingcomputationSFBSanta).
@@
A major upshot of the deep learning race is the emergence of spectacularly capable next-generation compute accelerators (zhang2016cambricon; emani2021accelerating; jia2019dissecting; medina2020habana). Although tailored expressly to deep learning workloads, these hardware platforms represent an exceptional opportunity to leapfrog progress on grand challenges in artificial life. The emerging class of fabric-based accelerators, led by the 850,000 core Cerebras CS-2 Wafer-Scale Engine (WSE) (lauterbach2021path; lie2022cerebras), holds particular promise as a vehicle for artificial life models. This architecture interfaces multitudinous processing elements (PEs) in a physical lattice, with PEs executing independently with private on-chip memory and interacting locally through a network-like interface.
In this work, we explore how such hardware might be recruited for large-scale digital evolution, demonstrating a genetic algorithm implementation tailored to the dataflow-oriented computing model of the CS-2 platform. Indeed, rapid advances in the capability of accelerator devices, driven in particular by market demand for deep learning operations, are anticipated to drive advances in agent-based model capabilities (perumalla2022computer). The upcoming CS-3 chip, for instance, supports clustering potentially thousands of constituent accelerators (moore2024cerebras).
@@
Unfortunately, scale complicates simulation observability. It is not uncommon for the volume and velocity of data streams from contemporary simulation to outstrip hardware bandwidth and storage capacity (osti_1770192). Extensive engineering effort will be required to ensure large-scale simulation retains utility in pursuing rigorous hypothesis-driven objectives.
Here, we confront just a single aspect of simulation observability within distributed evolutionary simulation: phylogenetic history (i.e., evolutionary ancestry relationships). Phylogenetic history plays a critical role in many evolution studies, across study domains and in vivo and in silico model systems alike (faithConservationEvaluationPhylogenetic1992; STAMATAKIS2005phylogenetics; frenchHostPhylogenyShapes2023; kim2006discovery; lewinsohnStatedependentEvolutionaryModels2023a; lenski2003evolutionary; moreno2021case). Phylogenetic analysis can trace the history of notable evolutionary events (e.g., extinctions, evolutionary innovations), but also characterize more general questions about the underlying mode and tempo of evolution (moreno2023toward; hernandez2022can; shahbandegan2022untangling; lewinsohnStatedependentEvolutionaryModels2023a). Particularly notable, recent work has used comparison of observed phylogenies against those produced under differing simulation conditions to test hypotheses describing underlying dynamics within real-life evo-epidemiological systems (giardina2017inference; voznica2022deep). Additionally, in silico, phylogenetic information can even serve as a mechanism to guide evolution in application-oriented domains (lalejini2024phylogeny; lalejini2024runtime; murphy2008simple; burke2003increased).
@@
Most existing artificial life work uses centralized tracking to maintain an exact, complete record of phylogenetic history comprising all parent-offspring relationships that have existed over the course of a simulation (ray1992evolution; bohm2017mabe; de2012deap; garwood2019revosim; godin2019apoget; dolson2024phylotrackpy). Typically, records of extinct lineages are pruned to prevent memory bloat (moreno2024analysis). Although direct tracking is well suited to serial simulation or centralized controller-worker schemes, runtime communication overheads and sensitivity to data loss impede scaling to highly distributed systems — particularly those with lean memory capacity like the Cerebras WSE (moreno2024analysis). To overcome this limitation, we have developed reconstruction-based approaches to in silico phylogenetic tracking (moreno2022hereditary). These approaches require no centralized data collection during simulation runtime; instead, they use post hoc comparisons among end-state agent genomes to deduce approximate phylogenetic history — akin to how DNA-based analyses describe natural history. Figure 1 summarizes this reconstruction-based strategy.
Although analogous work with natural biosequences is notoriously challenging and data-intensive (neyman1971molecular; lemmon2013high), the recently-developed hereditary stratigraphy annotation architecture is explicitly designed for fast, accurate, and data-lean reconstruction. Designed to attach on underlying replicators as a neutral annotation (akin to noncoding DNA), it is a general-purpose technique potentially applicable across diverse study domains (liben2008tracing; cohen1987computer; friggeri2014rumor). In a stroke of convergent thinking, ackley2023robust reports use of “bar code” annotations on his self-replicators to achieve a measure of coarse-grained lineage tracing.
@@
In this paper, we report new software and algorithms that harness the Cerebras Wafer-Scale Engine to enable radically scaled-up agent-based evolution while retaining key aspects of observability necessary to support hypothesis-driven computational experiments. Implementation comprises two primary aspects:
@@ Methods
steady retention policy
tilted retention policy
This section recounts the core mechanics of hereditary stratigraphy methods for phylogenetic tracking and describes new lightweight “surface” data structures used in this work. Then, we describe the asynchronous compute-communicate implementation strategy behind the testbed island-based genetic algorithm used for WSE validation and benchmarking experiments.
@@
In biology, mutational drift encodes ancestry information in DNA genomes. Our proposed methods are analogous, with ancestry information captured within genomes themselves rather than external tracking. Phylogenetic history can then be estimated after the fact, as Figure 1 depicts. This strategy reduces runtime communication and is resilient to germane modes of data loss (e.g., dropped messages, hardware crash-out). Note that, in addition to sampling end-state genomes, extracting “fossil” genomes over the course of simulation can allow lineages that end up going extinct to be included in phylogenetic reconstruction.
@@
Small differentia size intensifies need for a lean data structure to back differentia record management. Shrinking differentia to a single bit would be absurd if each is accompanied by a 32-bit generational timestamp. To delete timestamps, though, we need means to recalculate them on demand. As such, all described algorithms can deduce timestamps of retained differentia solely from their storage index and the count of elapsed generations.
Lastly, design of hstrat annotations must also consider how available storage space should be allocated across the span of history. In one possible strategy, which we call “steady” policy, retained time points would be distributed evenly across history. By contrast, under “tilted” policy more recent time points are preferred (han2005stream; zhao2005generalized). Note that prior hereditary stratigraphy work refers to them instead as “depth-proportional” and “recency-proportional resolution” (moreno2022hereditary). Comparisons of reconstruction quality have shown that tilted policy gives higher quality reconstructions from the same amount of reconstruction space in most — but not all — circumstances (moreno2024guide). This pattern follows an intuition that high absolute precision is more useful to resolving recent events than more ancient ones. In practice, it may be desirable to use a hybrid approach that allocates half of available annotation space to each strategy (moreno2024guide). The bottom panels of Figure 2 contrast steady versus tilted behaviors.
@@
At the outset of this project, several problematic aspects of porting existing hereditary stratigraphy algorithms to the WSE became apparent. Issues stemmed, in part, from a fundamental feature of these algorithms: organization of retained strata in contiguous, sorted order. This design imposes various drawbacks:
@@
Addressing these concerns required fundamental reformulation of hereditary stratigraphy conventions. To this end, we introduce a constant-time indexing scheme that maps each lineage’s fingerprint stream directly onto a fixed-width memory buffer. Differentia pruning occurs implicitly, as a result of resident differentia being overwritten by new placements. As before, care is taken to guarantee temporally representative collections of resident differentiae within the memory buffer over elapsed time. In this new “surface”-based approach, differentiae remain in the same buffer position until they are overwritten — as opposed to the existing “column”-based approach, where they are kept in sorted order. Figure 3 contrasts the two approaches.
@@
The Cerebras Wafer-Scale Engine comprises a networked grid of independently executing compute cores (Processing Elements or PEs). Each PE contains dedicated message-handling infrastructure, which routes tagged 32-bit packets (“wavelets”) to neighboring PEs and/or to be processed locally, according to a programmed rule set. Each PE contains 48kb of private on-chip memory wideth single clock cycle latency; communication to neighboring PEs, too, incurs low latency (buitrago2021neocortex). PEs support standard arithmetic and flow-control operations, as well as vectorized 32- and 16-bit integer and floating-point operations.
The WSE device is programmed by providing “kernel” code, written in the Cerebras Software Language (CSL), to be executed on each PE. This language’s programming model purposefully reflects characteristics of the underlying WSE architecture. CSL uses an event-driven framework, with tasks defined to respond to wavelet activation signals exchanged between — and within — PEs. Scheduling of active tasks occurs via hardware-level microthreading, which allows for some level of concurrency. Special faculties are provided for asynchronous send-receive operations that exchange buffered data between PEs.
@@
Our implementation unfolds according to a generational update cycle. Migration, depicted as blue-red arrows, is handled first. Each PE maintains independent immigration buffers and emigration buffers dedicated to each cardinal neighbor, depicted by Figure 4 in solid blue and red, respectively. On simulation startup, asynchronous receive operations are opened to accept genomes from each neighboring PE into its corresponding immigration buffer. At startup, additionally, each emigration buffer is populated with genomes copied from the population and an asynchronous send request is opened for each. Asynchronous operations are registered to on-completion callbacks that set a per-buffer “send-” or “receive-complete” flag variable. In this work, we size send buffers to hold one genome and receive buffers to hold four genomes. The main population buffer held 32 genomes.
Subsequently, the main update loop tests all completion flags. For each immigration flag that is set, buffered genomes are copied into the main population buffer, replacing randomly chosen population members. Then, the flag is reset and a new receive request is initiated. Likewise, for each emigration flag set, corresponding send buffers are re-populated with randomly sampled genomes from the main population buffer. Corresponding flags are then reset and new send requests are initiated. Figure 4(b) summarizes the interplay between send/receive requests, callbacks, flags, and buffers each generation (“main cycle”).
@@
We used a 96-bit genome for clade reconstruction trials, shown in Supplementary Figure 8 (moreno2024supplement). At the outset of simulation, the first 16 bits of founding genomes were randomized and, subsequently, were inherited without mutation, thus identifying descendants of the same founding ancestor. The next 80 bits were used for hereditary stratigraph annotation, 16 bits for a generation counter and the remaining 64 as a tilted-retention surface with single-bit differentiae. Other experiments use a 128-bit genome layout, with the first 32 bits used for a floating point “fitness” value and the generation counter upgraded from 16 to 32 bits.
@@
Python implementations of the surface tilted and steady algorithms both took around 4.2 microseconds per operation (standard error [SE] 50ns and 66ns; ). For context, this speed was about the measured time for a surface placement using a trivial calculation (SE 0.05; ). As expected, column implementations of steady and tilted fared much worse, taking about and the execution time per operation compared to the surface operations. In both cases, surface implementations significantly outperformed their column counterparts (Mann-Whitney U test, ).
Zig microbenchmarks clock tilted surface annotation updates at 230 ns per operation (SE 0.9ns; ). For context, this time is a little more than twice that required for a main memory access in contemporary computing hardware (markus2022memory). Our results measure the operation at the measured time for a trivial placement calculation (SE 0.03; ). Zig steady implementation clocks 511 ns per operation (SE 4; ), trivial (SE 0.8ns; ). Note that speedup of Zig implementations relative to Python reflects an intrinsic performance penalty due to interpreter overhead of Python evaluation. However, comparisons among Python implementations and among Zig implementations are nonetheless informative because all are on equal footing in this regard.
@@
Multiplied out to a full day, 17,000 generations per second turnover would elapse around 1.5 billion generations. With 32 individuals hosted per each of 850,000 PEs, the net population size would sit around 27 million at full CS-2 wafer scale. (Note, though, that available on-chip memory could support thousands of our very simple agents per PE, raising the potential for a net population size on the order of a billion agents.) A naive extrapolation estimates on the order of a quadrillion agent replications could be achieved hourly at full wafer scale for such a very simple model. We look forward to conducting more thorough benchmarking and scaling experiments in future work.
@@
These timings measure phylogeny tracking as approximately equivalent to that of the other aspects of simulation, combined. Given the highly minimalistic nature of the agent model and selection process, this result is highly promising. In actual use, most experiments will likely involve a much more sophisticated agent model, so relative overhead of tracking will be diminished. Additionally, these results do not include caching and coarsening strategies discussed above, which would speed tracking up considerably.
@@
To assess overall correctness of our Cerebras Software Language (CSL) surface algorithm implementation, we reviewed the clade structure of sample phylogenetic reconstructions from emulated WSE hardware. Full reconstruction quality testing of the underlying new surface algorithms themselves is provided in other recent work (moreno2024guide). (There, we found quality to be comparable to existing column algorithms.) For these experiments, we tagged genomes with a 16-bit randomly generated identifier at simulation startup, as shown in Supplementary Figure 8 (moreno2024supplement). We instantiated populations over a PE collective for durations of 25, 50, 100, and 250 generation cycles with neutral selection. Phylogenetic reconstruction used one end-state genome sampled per PE.
Figure 6 shows reconstructed phylogenies from each duration. As expected, taxa belonging to the same founding clade (shown by color) are reconstructed as more closely related to each other than to other taxa. Additionally, and also as expected, the number of distinct remaining founding clades diminishes monotonically with increasing simulation duration. These consistencies corroborate the general integrity of our CSL surface implementation. Note, however, the incidence of moderate overestimations for relatedness between independently-tagged clades throughout. As mentioned earlier, and discussed in greater depth elsewhere (moreno2024guide), this is an expected artifact of hereditary stratigraphy with single-bit differentiae. Applications requiring greater reconstruction precision can opt for larger differentia sizes and/or higher differentia counts.
@@
Finally, we set out to assess the performance of our pipeline at full wafer scale. For these experiments, we used the four-word, tracking-enabled genome layout, with the full first 32 bits containing a floating point fitness value. We defined two treatments: purifying-only conditions, where 33% of agent replication events decreased fitness by a normally-distributed amount, and adaption-enabled conditions, which added beneficial mutations that increased fitness by a normally-distributed amount, occurring with 0.3% probability. These beneficial mutations introduced the possibility for selective sweeps. As before, we used tournament size 5 for both treatments. We performed four on-hardware replicates of each treatment instantiated on 10k (), 250k () and 562.5k () PE arrays. We halted each PE after it elapsed 1 million generations.